# Effects of model misspecification on small area estimators

### v1

@@ Abstract
Nested error regression models are commonly used to incorporate observational unit specific auxiliary variables to improve small area estimates. When the mean structure of this model is misspecified, there is generally an increase in the mean square prediction error (MSPE) of Empirical Best Linear Unbiased Predictors (EBLUP). Observed Best Prediction (OBP) method has been proposed with the intent to improve on the MSPE over EBLUP. We conduct a Monte Carlo simulation experiment to understand the effect of mispsecification of mean structures on different small area estimators. Our simulation results lead to an unexpected result that OBP may perform very poorly when observational unit level auxiliary variables are used and that OBP can be improved significantly when population means of those auxiliary variables (area level auxiliary variables) are used in the nested error regression model or when a corresponding area level model is used. Our simulation also indicates that the MSPE of OBP in an increasing function of the difference between the sample and population means of the auxiliary variables.
Keywords: Model misspecification; Small area estimation; OBP; EBLUP
@@ 1 Introduction
In this section, for simplicity of exposition, we assume that a simple random sample (SRS) is selected for each small area. Suppose that the values of auxiliary variables are known for every unit in the finite population. Battese et al. (1988) introduced the following nested error regression model that links the response variable to the auxiliary variables for each unit of the finite population:
$y_{ij}=\bm{x}_{ij}^{\prime}\bm{\beta}+v_{i}+e_{ij},\quad i=1,\dots,m,\quad j=1,\dots,N_{i},$ (1.1)
@@
where $m$ is the number of areas; $N_{i}$ is known population size of area $i$; $v_{i}$’s are area-specific random effect; $e_{ij}$’s are sampling errors. We assume that $\{v_{i},\;i=1,\cdots,m\}$ and $\{e_{ij},\;i=1,\cdots,m;j=1,\cdots,N_{i}\}$ are independent with $v_{i}\sim N(0,\sigma^{2}_{v})$ and $e_{ij}\sim N(0,\sigma^{2}_{e})$. In general, regression coefficients $\bm{\beta}$ and the variance components $\sigma^{2}_{v}$ and $\sigma^{2}_{e}$ are unknown.
The following special case of the nested error regression model, often referred to as unit-context model, has received considerable attention in recent years; see, e.g., Newhouse et al. (2022). The model can be written as:
$y_{ij}=\bar{\bm{X}}_{i}^{\prime}\bm{\beta}+v_{i}+e_{ij},\quad i=1,\dots,m,\quad j=1,\dots,N_{i},$ (1.2)
where $\bar{\bm{X}}_{i}$ is a vector of known area-specific auxiliary variables. We make the same assumptions on the random effects $v_{i}$ and sampling errors $e_{ij}.$ These models offer an alternative approach when unit level auxiliary variables $x_{ij}$ (e.g., say from census records) are either unavailable or too outdated to be used as unit-specific auxiliary variables but area specific auxiliary variables are available (e.g., from geo-spatial data).
@@
Area-level model, introduced by Fay III and Herriot (1979), can be another possibility to handle lack of good unit-specific auxiliary variables. It models area level direct estimates of the parameter of interest to area-level auxiliary variables, and assumes the form:
$\bar{y}_{i}=\bar{\bm{X}}_{i}^{\prime}\bm{\beta}+v_{i}+\bar{e}_{i},\quad i=1,\dots,m,$ (1.3)
@@
where $\bar{y}_{i}$ is direct estimate from survey; $\bar{\bm{X}}_{i}$ is a vector of known area-level auxiliary variables; $v_{i}$’s are area-specific random effect; $\bar{e}_{i}$’s are sampling errors. It is assumed that $v_{i}$’s and $\bar{e}_{i}$’s are independent with $v_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma^{2}_{v})$, $\bar{e}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma^{2}_{e}/n_{i})$. The variance of sampling errors, $\sigma^{2}_{e}/n_{i}$, is commonly assumed to be known, but in practice they are estimated using a smoothing technique such as the one given in Otto and Bell (1995).
@@
The Best Predictor (BP) of a mixed effect in a linear mixed model is simply the conditional mean of the mixed effect given the data. Conceptually this BP can be obtained under a full specification of the distributions of the random effects and errors. When conditional mean of the random effects given the data is assumed to be a linear function of the sample observations, an explicit formula for BP can be obtained. The Best Linear Unbiased Predictor (BLUP) is obtained from the Best Predictor (BP) under the assumed linear mixed model when the unknown regression coefficients in the BP is replaced by the weighted least square estimators. Estimated BLUP (EBLUP) is obtained when the unknown variance components of the model are replaced by standard estimators (e.g., REML). The BLUP and EBLUP can be viewed as an estimated BP (EBP) of the mixed effects.
@@
Jiang et al. (2011) proposed Observed Best Predictors (OBP) for small area means under a linear mixed model in an attempt to reduce the effects of misspecification of the mean structure in the assumed model. We note that their OBP can be motivated as EBP when best predictive estimators (BPE) of model parameters, which minimize the observed mean squared prediction error (MPSE), are used in place of standard model parameter estimators. In the context of area level models, Jiang et al. (2011) used both theoretical and empirical studies to demonstrate that OBP outperforms EBLUP in terms of MPSE when the underlying linear mixed model is misspecified. Subsequently, Jiang et al. (2015) considered OBP for nested error regression model, where both the mean function and variance components are misspecified. Their simulations indicated that OBP may perform significantly better than EBLUP in terms of both overall MSPE and area-specific MSPE.
@@
In this paper, we investigate the effects of misspecified mean function and variance components on the predictive performances of existing small area estimators. The rest of the paper is organized as follows. In Section 2, we present numerical studies and the evaluation results in terms of both overall MSPE and area-specific MSPE. In Section 3, we summarize and conclude our paper by giving some practical guidelines on implementing OBP to handle model misspecification issues.
@@ 2 Simulations
The derivations of OBP procedure for both area-level and unit-level models can be found in the supplementary materials of Jiang et al. (2011). Also, since unit-context model is a special type of unit-level models, we can directly derive OBP by replacing $\bm{x}_{ij}$ by $\bar{\bm{X}}_{i}$.
@@
And, the corresponding area-level model is given by:
@@
    OBP under the assumed unit-context model (OBP-UC),
@@
    OBP under the basic Fay-Herriot (area-level) model with known sampling variance calculated from the simulated population of the small areas (OBP-FH),
@@
    OBP under the assumed unit-level model (OBP-UNIT),
@@
    EBLUP under the assumed unit-level model (EBLUP).
@@
Each scenario is independently simulated $K=1000$ times. The performance of the estimators (A)-(E), under the above simulation setups, are assessed in terms of both overall and area-specific MSPEs. The area-specific MSPE is defined as $\text{MSPE}(\hat{\bar{Y}}_{i})=\text{E}(\hat{\bar{Y}}_{i}-\bar{Y}_{i})^{2}$, where $\bar{Y}_{i}=N_{i}^{-1}\sum_{i=1}^{N_{i}}y_{ij}$ is the true small area mean, and $\hat{\bar{Y}}_{i}$ is the predicted value for $i$th area, either by OBP or EBLUP. In Monte Carlo simulations, MSPE for area $i$ and overall MSPE are approximated by:
@@
${\text{MSPE}}\approx\frac{1}{K}\sum_{i=1}^{K}\left[\frac{1}{m}\sum_{k=1}^{m}(\hat{\bar{Y}}_{i}^{(k)}-\bar{Y}_{i}^{(k)})^{2}\right],$ (2.6)
@@
Table 2.1: simulated overall MSPE under the heteroscedastic NER model (2.4)
$(m,b)$ DIRECT OBP-UC OBP-FH OBP-UNIT EBLUP
(40, 5) 1.475 0.689 0.630 1.775 1.474
(100, 5) 1.488 0.638 0.588 1.409 1.495
(400, 5) 1.493 0.614 0.567 1.352 1.508
(40, 10) 1.475 0.696 0.636 3.919 1.522
(100, 10) 1.488 0.646 0.595 2.364 1.513
(400, 10) 1.493 0.621 0.574 1.711 1.510
Table 2.1 reports the simulated overall MSPE for the various simulation conditions and estimators under the true underlying model (2.4). First of all, direct estimator is not affected by model misspecification because it is not model based. EBLUP behavior is similar to direct because it automatically assigns more weight to the sample mean when the model is weak. Surprisingly, the performance of OBP-UNIT is worse than OBP-UC in all cases considered. Moreover, note that the simulated overall MSPE is even larger than those of direct estimators and EBLUPs with $b=10$. Also, OBP-UC performs much better than OBP-UNIT. The performance of OBP-UC is similar to OBP-FH in which $\sigma^{2}_{ei}$ is calculated from the entire population. The findings suggest that in instances of significant model misspecification, OBP-UNIT may not be recommended.
As for the area-specific MSPEs, we utilize boxplots to display the distributions of the area-specific MSPEs associated with all the estimators. See Figure 2.1. The boxplot of OBP-UNIT shows much larger median MSPE and variability than those of OBP-UC and OBP-FH and in some cases worse than direct and EBLUP. OBP-UC shows slightly larger variability than OBP-FH. It might be because that in contrast to area-level models, unit-context models incorporate the uncertainty resulting from the estimation of model parameters, such as sampling variances.
Figure 2.1: Distributions of area-specific simulated MSPEs under model (2.4)
@@
The results based on $K=1000$ simulations are displayed in Table 2.2. As expected, in this case EBLUP now performs much better than the direct estimator as the assumed model is closer to the true model. OBP-UNIT performs better than direct estimator for larger $m$, but performs worse than EBLUP in all situations. The performance of OBP-UC and OBP-FH is similar, especially both of them perform much better than OBP-UNIT and performs better than EBLUP for large $m$,
Table 2.2: Overall simulated MSPE under the heteroscedastic NER model (2.7)
$m$ DIRECT OBP-UC OBP-FH OBP-UNIT EBLUP
40 18.243 1.798 1.837 18.230 1.532
100 18.339 1.341 1.385 7.958 1.514
400 18.422 1.085 1.099 3.122 1.509
2.1 Remark
To further investigate the failure of OBP with misspecified unit-specific auxiliary variables, we undertake a numerical study to evaluate the association between $|\bar{X}_{i}-\bar{x}_{i}|$ (absolute difference between the $i$th area sample mean and population mean) and $i$th area ${\text{MSPE}_{i}}$, approximated by Monte Carlo. First, we generate the finite population for $y$ using the following NER superpopulation model:
@@ 2.1 Remark
The generative processes for the finite population for $x$, $v_{i}$, and $e_{ij}$ remain consistent with the aforementioned model. Subsequently, we draw SRS samples of size $n_{i}=4$ from the population of each small areas.
In a simulation setting, when the number of units within a specific area is relatively small, it is often impossible to have sample mean $\bar{x}_{i}$ equal or even close to population mean $\bar{X}_{i}$. Therefore, in this subsection, we manually change the realizations of sample means by letting $\bar{x}_{i}=\bar{X}_{i}+\text{bias}$, where the bias term assumes negative, positive, and zero values. Figure 2.2 demonstrates that simulated ${\text{MSPE}_{i}}$ for the OBP with unit level auxiliary variables increases with the difference between $\bar{X}_{i}$ and $\bar{x}_{i}$ for area $i=1$ (for other areas, figures are similar). Notably, when $\bar{X}_{i}-\bar{x}_{i}=0$, i.e., for the unit-context model, simulated ${\text{MSPE}_{i}}$ is observed to be smaller than corresponding values for any other instances of $\bar{X}_{i}-\bar{x}_{i}$.

### v2

@@ Abstract
Nested error regression models are commonly used to incorporate unit specific auxiliary variables to improve small area estimates. When the mean structure of the model is misspecified, the design-based mean squared prediction error (MSPE) of Empirical Best Linear Unbiased Predictors (EBLUP) generally increases. The Observed Best Prediction (OBP) method has been proposed with the intent to improve on the design-based MSPE over EBLUP. In this paper, we conduct a Monte Carlo simulation experiments to understand the effect of misspsecification of mean structures on different small area estimators. Our findings suggest that the OBP using unit-level auxiliary variables does not outperform the EBLUP in terms of design-based MSPE, unless the number of small areas $m$ is extremely large. Conversely, the performance of OBP significantly improves when area-level auxiliary variables are employed. This paper includes both analytical and numerical evidence to demonstrate these observations, providing practical insights for addressing model misspecification in SAE.
Keywords: Small area predictors; OBP; EBLUP; MSPE
@@ 1 Introduction
Battese et al. (1988) introduced a nested error regression (NER) model that links the response variable to the auxiliary variables at the unit level:
$y_{ij}=\bm{x}_{ij}^{\prime}\bm{\beta}+v_{i}+e_{ij},\quad i=1,\cdots,m,\quad j=1,\cdots,N_{i},$ (1.1)
@@
where $m$ is the number of areas; $N_{i}$ is known population size of area $i$; $v_{i}$’s are area-specific random effect; $e_{ij}$’s are sampling errors. We assume that $\{v_{i},\;i=1,\cdots,m\}$ and $\{e_{ij},\;i=1,\cdots,m;j=1,\cdots,N_{i}\}$ are independent with $v_{i}\sim N(0,\sigma^{2}_{v})$ and $e_{ij}\sim N(0,\sigma^{2}_{e})$. The regression coefficients $\bm{\beta}$ and the variance components $\sigma^{2}_{v}$ and $\sigma^{2}_{e}$ are generally unknown. In this section, for simplicity of exposition, we assume that a simple random sample (SRS) of size $n_{i}$ is selected for small area $i$.
In many applications, unit-level auxiliary variables are not available or are outdated. In such cases, models like the following special case of the NER model (Newhouse et al., 2022), often referred to as the unit-context model, are used:
$y_{ij}=\bar{\bm{X}}_{i}^{\prime}\bm{\beta}+v_{i}+e_{ij},\quad i=1,\cdots,m,\quad j=1,\cdots,N_{i},$ (1.2)
where $\bar{\bm{X}}_{i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}\bm{x}_{ij}$ is a vector of known population means of auxiliary variables for area $i$. The same assumptions on the random effects $v_{i}$ and sampling errors $e_{ij}$ apply.
@@
Area-level model, introduced by Fay and Herriot (1979), can be another possibility to handle lack of good unit-specific auxiliary variables. An area level model that corresponds to the unit level model (1.1) is given by:
$\bar{y}_{i}=\bar{\bm{X}}_{i}^{\prime}\bm{\beta}+v_{i}+\bar{e}_{i},\quad i=1,\cdots,m,$ (1.3)
@@
where $\bar{y}_{i}$ is the unweighted sample mean; $v_{i}$’s are area-specific random effect; $\bar{e}_{i}$’s are sampling errors. It is assumed that $v_{i}$’s and $\bar{e}_{i}$’s are independent with $v_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma^{2}_{v})$, $\bar{e}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,D_{i})$, where $D_{i}$ is known sampling variance of $\bar{y}_{i}$. In practice, $D_{i}$’s are estimated using a smoothing technique such as the one given in Otto and Bell (1995). In this case, one may propose a smoothed estimator of $D_{i}$ as $s^{2}/n_{i}$, where $s^{2}$ is the pooled sample variance using data from all $m$ areas.
@@
The Best Predictor (BP) of a mixed effect in a linear mixed model is simply the conditional mean of the mixed effect given the data. The explicit formula for the BP of small area mean can be explicitly obtained for both unit level and area level models. As in Jiang et al. (2015), we assumed normality of both model and sampling errors. However, note that normality is not required for obtaining the BP of small area means. For example, the BP can be obtained under the assumption that the conditional means of the random effects given the data is a linear function of the sample observations; see, e.g., Ghosh and Lahiri (1987). The Best Linear Unbiased Predictor (BLUP) is obtained from the Best Predictor (BP) under the assumed linear mixed model when the unknown regression coefficients in the BP is replaced by the weighted least square estimators. Estimated BLUP (EBLUP) is obtained when the unknown variance components of the model are replaced by standard estimators (e.g., REML). The BLUP and EBLUP can be viewed as an estimated BP (EBP) of the mixed effects.
@@
Jiang et al. (2011) proposed Observed Best Predictors (OBP) for small area means under a linear mixed model in an attempt to reduce the effects of misspecification of the mean structure in the assumed model. We note that their OBP can be motivated as EBP when best predictive estimators (BPE) of model parameters, which minimize the observed mean squared prediction error, are used in place of standard model parameter estimators. In the context of area level models, Jiang et al. (2011) used both theoretical and empirical studies to demonstrate that OBP outperforms EBLUP in terms of design-based MPSE when the underlying linear mixed model is misspecified. Subsequently, Jiang et al. (2015) considered OBP for a nested error regression model, where both the mean function and variance components are misspecified. Their simulations indicated that OBP may perform significantly better than EBLUP in terms of both overall and area-specific design-based MSPE. However, our simulation studies indicate that OBP using unit-level auxiliary variables (OBP-UNIT) does not outperform EBLUP unless the number of areas $m$ is extremely large, which is unusual in small area estimation. We found that using area-level auxiliary variables, such as in the unit-context or area-level models, one can improve the performance of OBP.
@@
In this paper, we investigate the effects of misspecified mean function and variance components on the predictive performances of existing small area estimators. We also provide both analytical and numerical evidence explaining why OBP-UNIT may underperform in terms of design-based MSPE, and how the use of area-level auxiliary variables can enhance its effectiveness. The rest of the paper is organized as follows. In Section 2, we provide further details on OBP methodology and the reasoning why the OBP with area-level auxiliary variables outperforms the OBP with unit level auxiliary variables. In Section 3 we present numerical studies and the evaluation results in terms of both overall and area-specific design-based MSPE. Section 4 offers conclusions and practical guidelines for implementing OBP in the presence of model misspecification.
2 Analytical comparison between OBP with unit-level auxiliary variables and OBP with area-level auxiliary variables
Let $\bm{\psi}=(\bm{\beta}^{\prime},\sigma^{2}_{u},\sigma^{2}_{e})$ represent the vector of model parameters of the nested error regression model, and let $\tilde{\bm{\theta}}(\bm{\psi})=[\tilde{\theta}_{i}(\bm{\psi})]_{1\leq i\leq m}$ denote the vector of BPs for the true small area means $\bm{\theta}=[\theta_{i}]_{1\leq i\leq m}$. The design-based MSPE of $\tilde{\bm{\theta}}(\bm{\psi})$ is defined as
$\text{MSPE}(\tilde{\bm{\theta}}(\bm{\psi}))=\text{E}(|\tilde{\bm{\theta}}(\bm{\psi})-\bm{\theta}|^{2})=\sum_{i=1}^{m}\text{E}(\tilde{\theta}_{i}(\bm{\psi})-\theta_{i})^{2},$ (2.1)
where the expectation is with respect to the sample design. Following Jiang et al. (2011) and Jiang et al. (2015), the MSPE in (2.1) has an alternative expression, which is a key idea of the OBP. Namely, it leads to the fundamental equation of the OBP,
$\text{MSPE}(\tilde{\bm{\theta}}(\bm{\psi}))=\text{E}\{Q(\bm{\psi})\},$ (2.2)
and $Q(\cdot)$ is called the observed MSPE function. Under the nested error regression model, the $Q$ function can be expressed as:
$\begin{split}Q=&\bm{\beta}^{\prime}(\bar{\bm{X}}-\bm{G}\bar{\bm{x}})^{\prime}(\bar{\bm{X}}-\bm{G}\bar{\bm{x}})\bm{\beta}-2\bar{\bm{y}}^{\prime}\{(\bm{I}_{m}-2\bm{G})\bar{\bm{X}}+\bm{G}^{2}\bar{\bm{x}}\}\bm{\beta}\\ &+\bar{\bm{y}}^{\prime}\bm{G}^{2}\bar{\bm{y}}+\bm{1}_{m}^{\prime}(\bm{I}_{m}-2\bm{G})\hat{\bm{\mu}}^{2},\\ \end{split}$ (2.3)
where $\bar{\bm{X}}=(\bar{\bm{X}}_{i}^{\prime})_{1\leq i\leq m}$, $\bar{\bm{x}}=(\bar{\bm{x}}_{i}^{\prime})_{1\leq i\leq m}$, $\bar{\bm{x}}_{i}$ being the vector of unweighted sample means of the auxiliary variables for area $i$, $\bar{\bm{y}}=(\bar{y}_{i})_{1\leq i\leq m}$, $\bm{G}=\text{diag}\{n_{i}/N_{i}+(1-n_{i}/N_{i})n_{i}\sigma^{2}_{v}/(n_{i}\sigma^{2}_{v}+\sigma^{2}_{e}),1\leq i\leq m\}$ and $\hat{\bm{\mu}}^{2}=(\hat{\mu}^{2}_{i})_{1\leq i\leq m}$is the design-unbiased estimator of $(\bar{Y}^{2}_{i})_{1\leq i\leq m}$. The BPE of $\bm{\psi}$, denoted as $\hat{\bm{\psi}}$, is the minimizer of $Q(\bm{\psi})$ with respect to $\bm{\psi}$.
To understand why OBP performs optimally when area-level auxiliary variables are used, it is instructive to consider the case when all the unit-level model parameters $\bm{\psi}$ are known. Using the facts that $\text{E}(\bar{\bm{x}}-\bar{\bm{X}})=0$ and $\text{E}\{\bm{\bar{y}}^{\prime}\bm{G}^{2}(\bm{\bar{x}}-\bm{\bar{X}})\bm{\beta}\}=\text{cov}(\bar{\bm{y}}^{\prime}\bm{G}^{2},\bm{\beta}^{\prime}\bar{\bm{x}}^{\prime})$, we get
$\begin{split}\text{E}Q=&\text{E}\{\bm{\beta}^{\prime}(\bm{\bar{x}}-\bm{\bar{X}})^{\prime}\bm{G}^{2}(\bm{\bar{x}}-\bm{\bar{X}})\bm{\beta}\}+\cdots\end{split}$ (2.4)
where $\cdots$ are the terms involving finite population parameters and known model parameters. Since $\bm{G}^{2}$ is positive definite, the expression is minimized at $\bar{\bm{x}}-\bar{\bm{X}}=0$, for any fixed $\bm{\beta}$, indicating that the use of area-level auxiliary variables leads to optimal OBP performance.
@@ 3 Simulations
The more detailed derivations of OBP procedure for both area-level and unit-level models can be found in the supplementary materials of Jiang et al. (2011). Also, since unit-context model is a special type of unit-level models, we can directly derive OBP under a unit-context model by replacing $\bm{x}_{ij}$ by $\bar{\bm{X}}_{i}$.
@@
And the corresponding area-level model is given by:
@@
    OBP under the assumed unit-context model (3.2) (OBP-UC),
@@
    OBP under the basic Fay-Herriot (area-level) model (3.3) with smoothed direct variance estimates: $\hat{D}_{i}=s^{2}/n_{i}$, $s^{2}=(n-1)^{-1}\sum_{i=1}^{m}\sum_{i=1}^{n_{i}}(y_{ij}-\bar{y})^{2}$ and $n=\sum_{i=1}^{m}n_{i}$,
@@
    OBP under the assumed unit-level model(3.1) (OBP-UNIT),
@@
    EBLUP under the assumed unit-level model (3.1) (EBLUP-UNIT),
  • (F)
    EBLUP under the assumed unit-context model (3.2) (EBLUP-UC).
@@
Each scenario is independently simulated $K=1000$ times. The performance of the estimators (A)-(F), under the above simulation setups, are assessed in terms of both overall and area-specific design-based MSPEs. The area-specific design-based MSPE is defined as $\text{MSPE}(\hat{\bar{Y}}_{i})=\text{E}(\hat{\bar{Y}}_{i}-\bar{Y}_{i})^{2}$, where $\bar{Y}_{i}=N_{i}^{-1}\sum_{i=1}^{N_{i}}y_{ij}$ is the true small area mean, and $\hat{\bar{Y}}_{i}$ is the predicted value for $i$th area, either by OBP or EBLUP and E is with respect to the sample design. In Monte Carlo simulations, MSPE for area $i$ and overall MSPE are approximated by:
@@
${\text{MSPE}}\approx\frac{1}{m}\sum_{i=1}^{m}\left[\frac{1}{K}\sum_{k=1}^{K}(\hat{\bar{Y}}_{i}^{(k)}-\bar{Y}_{i}^{(k)})^{2}\right],$ (3.6)
@@
Table 3.1: Overall simulated design-based MSPE when the finite population is generated from the heteroscedastic NER model (3.4)
$(m,b)$ DIRECT OBP-UC OBP-FH OBP-UNIT EBLUP-UNIT EBLUP-UC
(40, 5) 1.475 0.689 0.698 1.775 1.474 0.687
(100, 5) 1.488 0.638 0.657 1.409 1.495 0.638
(400, 5) 1.493 0.614 0.637 1.352 1.508 0.613
(40, 10) 1.475 0.696 0.705 3.919 1.522 0.694
(100, 10) 1.488 0.646 0.664 2.364 1.513 0.645
(400, 10) 1.493 0.621 0.644 1.711 1.510 0.621
Table 3.1 reports the simulated overall design-based MSPE for the various simulation conditions and estimators when the finite population is generated based on the true underlying model (3.4). The direct estimator is unaffected by model misspecification as it is not model-based. EBLUP-UNIT behavior is similar to direct because it automatically assigns more weight to the sample mean when the model is weak. While OBP-UNIT was designed to handle model misspecification, it surprisingly performs poorly, even worse than the direct estimator and EBLUP-UNIT when $b=10$. OBP-UC and EBLUP-UC show similar and better performance than EBLUP-UNIT. The reason why EBLUP-UC also performs well could be that, given the large population size $N_{i}$ and the fact that the population values come from a common distribution across all areas, the population means $\bar{X}_{i}$ are roughly equal across areas. As a result, the unit-context model and the area-level model approximate the true model closely. Consequently, predictors using area-level auxiliary variables tend to outperform those using unit-level auxiliary variables. Overall, our findings suggest that in instances of significant model misspecification, OBP-UNIT may not be an effective choice.
As for the area-specific MSPEs, we utilize boxplots to display the distributions of the area-specific design-based MSPEs associated with all the estimators. See Figure 3.1. The boxplot of OBP-UNIT shows much larger median MSPE and variability than those of OBP-UC and OBP-FH and in some cases worse than direct and EBLUP-UNIT. OBP-UC shows slightly larger variability than OBP-FH. It might be because that in contrast to area-level models, unit-context models incorporate the uncertainty resulting from the estimation of model parameters, such as sampling variances.
Figure 3.1: Distributions of area-specific simulated design-based MSPEs when the finite population is generated from the heteroscedastic NER model (3.4)
@@
The results based on $K=1000$ simulations are displayed in Table 3.2. As expected, in this case EBLUP-UNIT outperforms the direct estimator since the assumed model is closer to the true model and EBLUP can borrow strengths from other related areas. While OBP-UNIT was expected to perform comparably to EBLUP-UNIT in this scenario, our simulation results indicate that OBP-UNIT performs actually worse than EBLUP-UNIT in terms of the simulated MSPE. When $m=40$, the simulated MSPE for OBP-UNIT is similar to that of the DIRECT estimator, indicating poor performance. In contrast, both OBP-UC and EBLUP-UC show similar results, performing slightly better than EBLUP-UNIT when $m=100$ and $m=400$. OBP-FH demonstrates stable performance and performs better than OBP-UC and EBLUP-UC when $m=40$.
The results in Tables (3.1) and (3.2) indicates that OBP-UNIT may not effectively reduce the impact of model misspecification, compared with EBLUP-UNIT. This surprising outcome prompted further investigation, which we explore in the following Remarks 1 and 2.
Table 3.2: Overall simulated design-based MSPE when the finite population is generated from the heteroscedastic NER model (3.7)
$m$ DIRECT OBP-UC OBP-FH OBP-UNIT EBLUP-UNIT EBLUP-UC
40 18.243 1.798 1.588 18.230 1.532 1.872
100 18.339 1.341 1.248 7.958 1.514 1.358
400 18.422 1.085 1.059 3.122 1.509 1.087
Remark 1: MSPE of OBP-UNIT vs $|\bar{\bm{X}}-\bar{\bm{x}}|$
To further investigate the relationship between the performance of OBP-UNIT and the absolute difference between sample means and population means, we conducted a numerical study using the following Nested Error Regression (NER) superpopulation model:
@@
The generative processes for the finite population for $x$, $v_{i}$, and $e_{ij}$ remain consistent with the aforementioned model. Next, we drew SRS samples of size $n_{i}=4$ for each small areas.
In a simulation setting, when the number of units within a specific area is relatively small, it is often impossible to have sample mean $\bar{\bm{x}}$ equal or even close to population mean $\bar{\bm{X}}$. To explore how deviations of sample means $\bar{\bm{x}}$ from population means $\bar{\bm{X}}$ affect the performance of OBP, we introduced a bias term in the sample means. Specifically, for each area, the sample means were adjusted by setting: $\bar{\bm{x}}=\bar{\bm{X}}+\text{bias}$, where the bias term varied to take on negative, positive, and zero values.
The design-based MSPE for OBP using unit-level auxiliary variables ($\text{MSPE}_{i}$) was calculated for each area, and these values were compared with the magnitude of the absolute difference $|\bar{\bm{X}}-\bar{\bm{x}}|$. The results were visualized in Figure 3.2, which illustrates that the simulated ${\text{MSPE}_{i}}$ for OBP-UNIT increases with the difference between $\bar{X}_{i}$ and $\bar{x}_{i}$ for area $i=1$. Notably, when $\bar{X}_{i}-\bar{x}_{i}=0$, corresponding to the unit-context model, simulated ${\text{MSPE}_{i}}$ is lower compared to other scenarios with non-zero differences. These results align with our theoretical design-based expectation of $Q$ in section 2.